OpenStudy (anonymous):

Please help me with this problem! I have a bit of trouble with it. 1+sec x/tan x + sin x = csc x I get stuck halfway through.

OpenStudy (anonymous):

is that (1+secx)/tanx......

OpenStudy (anonymous):

It's (1 + secx)/(tanx+sinx)

OpenStudy (anonymous):

I have this written down so far.

OpenStudy (anonymous):

Try turning it into (1+secx)*(cotx+cscx) and foiling it out

OpenStudy (anonymous):

I don't get a square at the end. I end up getting cotx+cscx+secxcotx+cscxsecx

OpenStudy (anonymous):

unless you want me to simplify it?

OpenStudy (anonymous):

which cancels to cotx+cscx+cscx+1/sinxcosx

OpenStudy (anonymous):

cotx+2cscx+2/sin2x

OpenStudy (anonymous):

which dosent work... hymm let me try something else

OpenStudy (anonymous):

i know some cosines are supposed to cross out but i'm not very sure after that.

OpenStudy (anonymous):

unless the 1 should be converted to cosx/cosx?

OpenStudy (anonymous):

did you try dividing the right hand side into the left, leaving (....)=1 i know the tanx will cancel then

OpenStudy (anonymous):

then you get 1+secx/1+cosx = 1

OpenStudy (anonymous):

but the answer is cscx? hmmm

OpenStudy (anonymous):

well, what you have is an equation, i would steer clear of thinking one side is the "answer" of the other, it is algebraically allowed to divide the LHS (left hand side) by the RHS leaving (1+secx)/(1+cosx)=cscx/cscx

OpenStudy (anonymous):

ah that's true

OpenStudy (anonymous):

OOPS, i made a mistake

OpenStudy (anonymous):

i tried to solve it again and i get this so far

OpenStudy (anonymous):

ok so what you want to do is multiply the right hand side by (sinx+tanx)/(sinx+tanx)

OpenStudy (anonymous):

if you do that then you will see that they are equal

OpenStudy (anonymous):

wait the bottom can be factored...

OpenStudy (anonymous):

because sinxcosx+sinx also equals sinx(1+cosx)

OpenStudy (anonymous):

\[\frac{ 1+\sec x }{ \tan x+\sin x }=\csc x * (\frac{ \sin x+tanx }{ \sin x+tanx })\]

OpenStudy (anonymous):

^^^ this is correct

OpenStudy (anonymous):

let me try it out

OpenStudy (anonymous):

do you see what i did, the RHS is being multiplied by one

OpenStudy (anonymous):

rhs? sorry i'm not sure what that is

OpenStudy (anonymous):

Right hand side, the side that has the cscx term

OpenStudy (anonymous):

OH okay

OpenStudy (anonymous):

a question i wish i would have asked my self sooner on this problem is "can i multiply one of the sides by 1 (in some form) to make it equal the other"

OpenStudy (anonymous):

oh yeah, that's right

OpenStudy (anonymous):

i forgot about that

OpenStudy (anonymous):

so did i lol

OpenStudy (anonymous):

Here is what I tell my students to do. Look at the more complex side and turn everything into terms of sines and cosines.\[\frac{ 1+\sec \theta }{ \tan \theta+\sin \theta }=\frac{ 1+\frac{ 1 }{ \cos \theta } }{ \frac{ \sin \theta }{ \cos \theta } +\sin \theta}\] Then get a common denominator on top and on bottom\[=\frac{ \cos \theta+1 }{ \cos \theta } \div \frac{ \sin \theta+\sin \theta \cos \theta }{ \cos \theta }\] Then multiply by reciprocal and factor a sine theta\[=\frac{ \cos \theta+1 }{ \cos \theta } \times \frac{ \cos \theta }{ \sin \theta \left( 1+\cos \theta \right) }\] \[=\frac{ 1 }{ \sin \theta } \] \[=\csc \theta \]

OpenStudy (anonymous):

@goberds and @Edutopia

OpenStudy (anonymous):

I dont understand what you did on the "common denominator step"

OpenStudy (anonymous):

did you multiply both sides by the reciprocal of the RHS?

OpenStudy (anonymous):

1+1/cosx =(cosx+1)/cosx and sinx/cosx +sinx = (sinx +sinxcosx)/cosx

OpenStudy (anonymous):

\[1+\frac{ 1 }{ \cos \theta}=\frac{ \cos \theta }{ \cos \theta }*1+\frac{ 1 }{ \cos \theta }\] equals \[\frac{ \cos \theta+1 }{ \cos \theta } \]

OpenStudy (anonymous):

Then I did the same thing for the denominator.